# Federated Unsupervised Semantic Segmentation

### v1

@@ 3.1 Local training
Following the centralized USS training of (Hamilton et al., 2022), each client in FUSS is initialized with a frozen VFM encoder, enabling efficient and stable local learning. This choice aligns with recent trends in efficient FL adaptation (Wang et al., 2024), though FUSS takes a simpler approach: by freezing the backbone of the encoder entirely, each client performs one-time feature extraction over its local dataset, and optimizes and transmits only a segmentation head $S_{k}$ and the centroid matrix $\mathbf{M}_{k}$. This eliminates the need for layer scheduling or repeated backward passes, enabling streamlined training even in resource-constrained settings.
@@
$\mathbf{Z}_{k_{i}}=E_{\text{VFM}}(\mathbf{X}_{k_{i}},\theta_{\text{VFM}})\in\mathbb{R}^{H\times W\times D^{\prime}},$
@@
$\mathbf{H}_{k_{i}}=S_{k}(\mathbf{Z}_{k_{i}},\theta_{S_{t}})\in\mathbb{R}^{H\times W\times D}.$
To optimize the segmentation head, since $D>D^{\prime}$ FUSS, instead of a loss in the form of Eq. 2, adopts a correspondence distillation loss that encourages semantic consistency across images, across features projected in different dimensions. Given feature tensors $\mathbf{Z}_{k_{i}}$ and $\mathbf{Z}_{k_{j}}$ from two images, we define a cosine similarity tensor:
$\mathbf{A}_{k_{i,j}}^{(h,w,m,n)}:=s_{\cos}(\mathbf{Z}_{k_{i}}^{(h,w)},\mathbf{Z}_{k_{j}}^{(m,n)}),$
@@
$\mathbf{Q}_{k_{i,j}}^{(h,w,m,n)}:=s_{\cos}(\mathbf{H}_{k_{i}}^{(h,w)},\mathbf{H}_{k_{j}}^{(m,n)}).$
@@
Input : Clients $\mathcal{K}$, pretrained VFM $E_{\text{VFM}}$, projection head $S_{k}$, centroids $\mathbf{M}_{k}$, rounds $R$
Output : Global segmentation head $\bar{\theta}_{S}$ and centroids $\bar{\mathbf{M}}$
@@
             Initialize $\theta_{S_{t}}\leftarrow\bar{\theta}_{S}$, $\mathbf{M}_{k}\leftarrow\bar{\mathbf{M}}$
             for each local step do
@@
                   Compute $\mathcal{L}_{\text{corr}},\mathcal{L}_{\text{cluster}}$
                   Update $\theta_{S_{k}},\mathbf{M}_{k}$
@@
      Server aggregates into $(\bar{\theta}_{S},\bar{\mathbf{M}})$
       Server broadcasts parameters to all clients
@@ 3.2.1 Extended Federated Averaging
We adapt the canonical Federated Averaging algorithm (McMahan et al., 2017) to our setting by aggregating both the segmentation head and the centroid matrix. While standard FedAvg aggregates encoder weights, it does not account for latent class prototypes — a critical element in FUSS. Let $\theta_{S_{k}}$ denote the segmentation head parameters and $\mathbf{M}_{k}\in\mathbb{R}^{|\mathcal{C}|\times D}$ the centroid matrix from client $k$. The aggregated global parameters are:
@@ 3.2.2 Centroid Clustering for Prototype Alignment (FedCC)
We propose Federated Centroid Clustering (FedCC), a family of novel aggregation strategies that cluster the union of local prototypes across clients. Rather than assuming consistent object-to-centroid mappings across clients, FedCC clusters the pooled centroids globally—treating all received centroids as unordered semantic candidates. This strategy remains agnostic to local class semantics and promotes emergent alignment through global structure discovery in prototype space.
@@
$\mathbf{m}_{\text{pool}}=\bigcup_{k\in\mathcal{K}}\{\mathbf{m}_{k_{c}}\mid c=1,\dots,|\mathcal{C}|\}\in\mathbb{R}^{(|\mathcal{K}|\cdot|\mathcal{C}|)\times D}.$ (10)
@@
(a) FedCC - $k$Means Clustering.
@@
(b) FedCC - Maximin Prototype Selection.
@@
In both variants, $\bar{\theta}_{S}$ is updated via Eq. (8), and the resulting pair $(\bar{\theta}_{S},\bar{\mathbf{M}})$ is broadcast to all clients. The two variants of FedCC serve complementary goals: FedCC–$k$Means promotes semantic consensus by averaging structurally similar centroids across clients, while FedCC–Maximin promotes semantic diversity, selecting centroids that span the prototype space. The former is suitable when clients share overlapping classes with misaligned feature clusters, whereas the latter guards against prototype collapse class-imbalanced settings.
@@ 4.1 Experimental Setup
We conduct experiments on the following benchmarks: Cityscapes (Cordts et al., 2016), a 1/21 subsampled version of CocoStuff (Caesar et al., 2018), and a private industrial dataset referred to as IPS (Industrial Pipeline Segmentation).
@@
Cityscapes is used to simulate a cross-silo federation: we reserve all samples from three cities for validation, and partition the remaining cities among clients. Depending on the federation size (3, 6, or 18 clients), each client is assigned images from 6, 3, or 1 city respectively, promoting spatial and semantic locality in client data. More details can be found in Appendix C.
CocoStuff, while not naturally suited to real-world federated use cases, is included to test our method’s robustness under extreme data heterogeneity. We assign each image a dominant class based on pixel-wise frequency, and apply Dirichlet sampling to partition the dataset across clients, inducing strong class imbalance and inter-client variability. More details can be found in Appendix D.
To evaluate real-world applicability, we have employed the IPS (Psarras et al., 2024) dataset, which contains high-resolution RGB images and video frames captured at four industrial sites performing pipeline infrastructure inspection. IPS poses a binary semantic segmentation task with two classes: background and pipe, and contains 984 training samples. We simulate a 3-client federation by assigning each client to a distinct site, while a fourth, unseen site is reserved exclusively for validation—assessing generalization under cross-domain deployment. Visual examples can be found in Appendix E.
4.2 Implementation Details
Each client is equipped with a frozen Vision Foundation Model (VFM) backbone, specifically the base version of DINO (Caron et al., 2021) pretrained on ImageNet (Russakovsky et al., 2015), which is used to extract semantically rich features from raw input images. On top of this backbone, a trainable segmentation head $S(\cdot,\theta_{S})$ is implemented as a lightweight two-layer convolutional projection network that maps the VFM 768-dimensional pixel embeddings, to a lower-dimensional embedding space of 70 dimensions.
Training of each local segmentation head $S_{k}$ is performed using mini-batches of 8 query images sampled from the client’s dataset. For each query image, we select one nearest neighbor (by feature similarity) and five additional random support images from the same client. The cross-image correlation loss (Eq. 6) is computed between the query and each support image individually. We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $5\times 10^{-4}$ to optimize $\theta_{S_{k}}$.
Table 1: Experimental results comparing naive and FedCC versions of FedAvg, FedProx, and FedMoon across different client counts. Reported metric is mean Intersection-over-Union (mIoU). CRF post proccessing is applied to all results.
Method Cityscapes IPS
3 Clients 6 Clients 18 Clients 3 Clients
FedProx 19.77 20.01 18.22 84.63
FedProx+FedCC 20.96 20.48 20.90 83.52
FedMOON 19.83 19.94 19.86 83.78
FedMOON+FedCC 20.35 21.76 20.41 84.70
FedAvg 21.12 19.94 21.25 84.10
FedAvg+FedCC 21.83 21.80 21.53 84.87
Centralized 21.0 85.66
4.3 Results
To provide broader context, we also incorporate two common local training regularizers—FedProx (Li et al., 2020) and FedMoon (Li et al., 2021)—which are orthogonal to our aggregation strategies but useful for understanding FedCC’s modularity and robustness under both i.i.d. and non-i.i.d. conditions. More information about their implementation can be found at Appendix B.
Method mIoU
FedProx 22.39
FedMoon 24.69
FedAvg 24.69
FedProx + FedCC 24.38
FedMoon + FedCC 23.15
FedAvg + FedCC 25.39
Centralized 25.41
A qualitative assessment of the results under heterogeneous data distributions is presented in Figure 1. These examples illustrate that even in strongly non-i.i.d. scenarios, the centroids produced by FedCC remain structurally discriminative across the entire federation. Furthermore, visualizations of the projected centroid embedding space reveal that while FedAvg tends to produce cluttered prototypes, FedCC maintains significantly greater inter-class separability.
To isolate the contributions of each component in FUSS, we perform an ablation study on the aggregation process—evaluating the effects of weighted averaging (W), encoder parameter fusion (E), and prototype aggregation (C). Table 3 presents results across multiclass (Cityscapes) and binary (IPS) segmentation.
Table 3: Ablation study for multiclass (Cityscapes) and binary (IPS) federated unsupervised semantic segmentation. We evaluate the impact of three aggregation modules: W = weighted averaging (by client dataset size), E = encoder aggregation, C = centroid aggregation. Reported scores are mean Intersection-over-Union (mIoU). For configurations with no resulting global model, we report the average result along with the best and worst-performing clients.
Method Agg. Modules Cityscapes IPS
W E C 3 Clients 6 Clients 18 Clients 3 Clients
No Agg. $18.60_{-0.69}^{+0.72}$ $18.20^{+1.07}_{-1.62}$ $16.74^{+1.71}_{-1.62}$ $67.24^{+15.0}_{-27.0}$
FedAvg 18.55 +${}_{-0.7}^{+0.6}$ $17.97^{+0.72}_{-0.48}$ $17.18^{+1.52}_{-1.58}$ $70.74^{+12.4}_{-18.0}$
19.26 18.22 17.55 80.64
18.55 +${}_{-0.7}^{+0.7}$ $18.45^{+0.47}_{-1.22}$ $17.36^{+1.42}_{-2.20}$ $78.43^{+4.78}_{-4.73}$
19.24 18.96 17.50 79.16
[…]
18.88 19.41 17.89 80.67
19.71 19.39 17.87 82.60
[…]
19.50 19.38 18.12 78.22
19.52 19.64 18.87 80.27
The aggregation module analysis confirms the benefit of explicitly aligning prototypes (C), even without encoder averaging. For instance, C-only FedAvg outperforms E-only FedAvg across the majority of settings. On IPS, FedCC-KMeans performs best, likely due to strong foreground-background separability. However, for Cityscapes—where small objects and semantic clutter dominate—FedCC-Maximin proves more effective by promoting inter-centroid diversity and mitigating prototype collapse.
Limitations
5 Conclusion
We addressed a critical gap in existing federated learning literature by proposing a general formulation that integrates unsupervised representation learning with class prototype alignment across clients. Through a combination of correlation-based local optimization and novel aggregation strategies—including variations of weighted parameter averaging and global clustering procedures—we demonstrated that FUSS effectively mitigates challenges such, semantic misalignment, and local overfitting.
Extensive experiments on both benchmark (Cityscapes, CocoStuff) and real-world (IPS) datasets show that our method consistently outperforms local-only and traditional Federated Learning training baselines, achieving segmentation quality competitive with centralized unsupervised models, while remaining scalable and robust in i.i.d and non-i.i.d data regimes.
Future extensions may explore end-to-end learnable aggregation schemes, adaptive client weighting strategies based on local task uncertainty, or integration with semi-supervised constraints when limited labels become available.
Appendix A FUSS framework visualization
Figure 2: Overview of the proposed FUSS framework.
Figure 2 provides a visual overview of the proposed FUSS framework, as described in Section 3. Each client performs local unsupervised training using self-supervised objectives and periodically transmits its learned segmentation head and semantic prototypes to a central server. The server aggregates these parameters and broadcasts the updated global model back to all clients.
While traditional federated strategies may result in disorganized or overlapping class prototypes due to local heterogeneity, our proposed aggregation methods explicitly promote semantic alignment and prototype separability across the federation.
Appendix B Client-Side Regularization Baselines
FedProx. FedProx [Li et al., 2020] adds a proximal term to the local loss to prevent divergence from the global model under non-i.i.d. data. In FUSS, we apply it to the segmentation head:
$\mathcal{L}_{\text{FedProx}}=\mathcal{L}_{\text{local}}+\frac{\mu}{2}\left\|\theta_{S_{k}}-\bar{\theta}_{S}\right\|^{2},$
FedMoon. FedMoon [Li et al., 2021] introduces a contrastive loss to align local and global representations. In FUSS, we apply this to the segmentation embeddings:
$\mathcal{L}_{\text{Moon}}=-\log\left(\frac{\exp(s_{\cos}(\mathbf{z}_{k},\bar{\mathbf{z}})/\tau)}{\exp(s_{\cos}(\mathbf{z}_{k},\bar{\mathbf{z}})/\tau)+\exp(s_{\cos}(\mathbf{z}_{k},\mathbf{z}^{\text{prev}}_{k})/\tau)}\right),$
and define the final objective as $\mathcal{L}_{\text{FedMoon}}=\mathcal{L}_{\text{local}}+\lambda\cdot\mathcal{L}_{\text{Moon}}$. This promotes representation consistency across rounds.
Appendix C Cityscapes Federation Scenarios
@@
Table 4: Example of Cityscapes split for applied federation scenarios
@@
Each federation client $k$ receives a disjoint subset $\mathcal{D}_{k}$ composed of images from geographically distinct cities, introducing domain distribution shifts. An example of the Cityscape split for each federation simulation can be observed in Table 4. We also apply a five-crop on each image as a data augmentation technique, to better handle training times and small objects.
C.1 Cityscapes qualitative results
Appendix D CocoStuff non-i.i.d Sampling
We construct non-i.i.d. partitions of the COCO-Stuff dataset using a Dirichlet-based sampling strategy informed by semantic supercategories. The process involves assigning each image to a single coarse semantic class based on pixel frequency, followed by Dirichlet sampling to allocate images to clients in a skewed but controllable manner.

### v2

@@ 4.1 Local training
Following the centralized USS training of [14], each client in FUSS is initialized with a frozen VFM encoder, enabling efficient and stable local learning. This choice aligns with recent trends in efficient FL adaptation [40], though FUSS takes a simpler approach: by freezing the backbone of the encoder entirely, each client performs one-time feature extraction over its local dataset, and optimizes and transmits only a segmentation head $S_{k}$ and the centroid matrix $\mathbf{M}_{k}$. This eliminates the need for layer scheduling or repeated backward passes, enabling streamlined training even in resource-constrained settings.
@@
$\mathbf{Z}_{k_{i}}=E_{\text{VFM}}(\mathbf{X}_{k_{i}},\theta_{\text{VFM}})\in\mathbb{R}^{H\times W\times D^{\prime}},$ (6)
@@
$\mathbf{H}_{k_{i}}=S_{k}(\mathbf{Z}_{k_{i}},\theta_{S_{k}})\in\mathbb{R}^{H\times W\times D}.$ (7)
To optimize the segmentation head, since $D<<D^{\prime}$ FUSS, instead of a loss in the form of Eq. 2, adopts a correspondence distillation loss that encourages semantic consistency across images, across features projected in different dimensions. Given feature maps $\mathbf{Z}_{k_{i}}$ and $\mathbf{Z}_{k_{j}}$ from two images, we define a cosine similarity tensor:
$\mathbf{A}_{k_{i,j}}^{(h,w,m,n)}:=s_{\cos}(\mathbf{Z}_{k_{i}}^{(h,w)},\mathbf{Z}_{k_{j}}^{(m,n)}),$ (8)
@@
$\mathbf{Q}_{k_{i,j}}^{(h,w,m,n)}:=s_{\cos}(\mathbf{H}_{k_{i}}^{(h,w)},\mathbf{H}_{k_{j}}^{(m,n)}).$ (9)
@@
This architectural design offers crucial advantages for Federated Learning. First, by restricting optimization to a lightweight segmentation head, we drastically reduce the computational burden on clients and minimize the communication overhead, as only the small projection head $S_{k}$ and centroid matrix $\mathbf{M}_{k}$ need to be transmitted. Second, this formulation is mathematically ideal for the federated setting because it explicitly produces discrete cluster centroids. These centroids act as semantically rich prototypes that geometrically summarize the local data distribution of each client. In our proposed methodology, these prototypes serve as a “common language” or set of anchor points, enabling the server to align discovered semantic classes across heterogeneous clients in a robust and efficient manner.
Input : Clients $\mathcal{K}$, pretrained VFM $E_{\text{VFM}}$, projection head $S_{k}$, centroids $\mathbf{M}_{k}$, aggregation rounds $R$, correlation learning rate $\eta_{corr}$, centroid learning rate $\eta_{clust}$
[…]
@@
      Initialize $\theta_{S_{k}}\leftarrow\bar{\theta}_{S}$, $\mathbf{M}_{k}\leftarrow\bar{\mathbf{M}}$
    for each local step $i$ do
@@
       // Step 1: Compute Correlation loss
         Compute $\mathcal{L}_{\text{corr}}$ using Eq. (10)
       // Step 2: Compute Cluster Loss
         Compute $\mathcal{L}_{\text{cluster}}$ using detached features $\mathbf{H}_{k_{i}}$
         in Eq. (5)
       // Step 3: Update Parameters
       $\theta_{S_{k}}\leftarrow\theta_{S_{k}}-\eta_{corr}\nabla_{\theta_{S_{k}}}\mathcal{L}_{\text{corr}}$
       $\mathbf{M}_{k}\leftarrow\mathbf{M}_{k}-\eta_{clust}\nabla_{\mathbf{M}_{k}}\mathcal{L}_{\text{cluster}}$
@@
  Server aggregates $(\theta_{S_{k}},\mathbf{M}_{k}),\text{with }k=1,...,|\mathcal{K}|,$ into $(\bar{\theta}_{S},\bar{\mathbf{M}})$
   Server broadcasts $(\bar{\theta}_{S},\bar{\mathbf{M}})$ to all clients
@@ 4.2.1 Extended Federated Averaging
First we extend the canonical Federated Averaging algorithm [27] to our setting by aggregating both the segmentation head parameters and the centroid matrices. While standard FedAvg aggregates encoder weights, it does not account for latent class prototypes, a critical element in FUSS. Let $\theta_{S_{k}}$ denote the segmentation head parameters and $\mathbf{M}_{k}\in\mathbb{R}^{|\mathcal{C}|\times D}$ the centroid matrix from client $k$. The aggregated global parameters are:
@@ 4.2.2 Centroid Clustering for Prototype Alignment (FedCC)
We propose Federated Centroid Clustering (FedCC), a family of novel aggregation strategies that cluster the union of local prototypes across clients. Rather than assuming consistent object-to-centroid mappings across clients, FedCC clusters the pooled centroids globally by treating all received centroids as unordered semantic candidates. This strategy remains agnostic to local class semantics and promotes emergent alignment through global structure discovery in prototype space.
@@
$\mathcal{M}_{\text{pool}}=\bigcup_{k\in\mathcal{K}}\{\mathbf{m}_{k_{c}}\mid c=1,\dots,|\mathcal{C}|\}\in\mathbb{R}^{(|\mathcal{K}|\cdot|\mathcal{C}|)\times D}.$ (14)
@@
FedCC - kkMeans Clustering
@@
FedCC - Maximin Prototype Selection
@@
In both variants, $\bar{\theta}_{S}$ is updated via Eq. 12, and the resulting pair $(\bar{\theta}_{S},\bar{\mathbf{M}})$ is broadcast to all clients. The two variants of FedCC serve complementary goals: FedCC–kkMeans promotes semantic consensus by averaging structurally similar centroids across clients, while FedCC–Maximin promotes semantic diversity, selecting centroids that span the prototype space. The former is suitable when clients share overlapping classes with misaligned feature clusters, whereas the latter guards against prototype collapse class-imbalanced settings.
Figure 1 provides a visual overview of the proposed FUSS framework.
4.3 Theoretical Guarantees and Convergence Behavior
While FedAvg remains a cornerstone of many Federated Learning algorithms, even in modified forms, we question whether it can be made more accurate and reliable in situations where it underperforms. The FUSS setting poses a unique technical challenge: in the absence of labels or supervision, clients must align high-dimensional, semantically meaningful features across heterogeneous and potentially disjoint data distributions. Without supervision to guide inter-client alignment, naive averaging of learned class prototypes may fail, especially under non-i.i.d. data, by merging distinct classes or collapsing ambiguous regions.
While proving convergence in the fully unsupervised and federated setting remains a complex open problem currently being investigated by the research community, we offer the following theoretical justification based on known results for FedAvg and our prototype aggregation design.
Let
$\mathbf{M}_{i}=\big[\mathbf{m}_{i}^{(1)},\dots,\mathbf{m}_{i}^{|\mathcal{C}|}\big]\in\mathbb{R}^{|\mathcal{C}|\times d}$ (17)
denote the class prototype matrix (centroids) computed by client $i$, where $|\mathcal{C}|$ is the number of semantic clusters and $d$ the feature dimension. FedAvg-style aggregation computes the global prototype per class via:
$\bar{\mathbf{m}}^{(c)}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{m}_{i}^{(c)},$ (18)
which corresponds to the class $c$ average across $N$ clients.
Aligned Case
We first argue that in cases of correct semantic alignment and global prototype convergence through FedAvg, replacing centroid averaging with FedCC does not introduce instability and will also converge. Suppose that all clients produce consistent class prototypes, i.e.,
$\mathbf{m}_{1}^{(c)}\approx\mathbf{m}_{2}^{(c)}\approx\dots\approx\mathbf{m}_{N}^{(c)}.$ (19)
In this case, FedAvg and FedCC (which, in the case of Maximin, selects a maximally diverse prototype among clients) both yield similar and semantically consistent centroids:
$\mathbf{m}^{(c)}_{\text{FedCC}}\approx\bar{\mathbf{m}}^{(c)}.$ (20)
Misaligned Case
If clients produce misaligned or collapsed prototypes, averaging may produce ambiguous or meaningless centroids:
$\max_{i,j}\left\|\mathbf{m}_{i}^{(c)}-\mathbf{m}_{j}^{(c)}\right\|\gg 0$ (21)
that would lead $\bar{\mathbf{m}}^{(c)}$ to lie in-between distinct classes / be a collapsed representation. In contrast, FedCC-Maximin selects the prototype $\mathbf{m}^{(c)}_{\text{FedCC}}$ that maximizes its minimum distance to all previously selected ones as shown in Eq. 16, pushing towards greater inter-class diversity.
On the other hand, while FedCC-kMeans will still be problematic due to the averaging nature of the algorithm, will at least find structurally consistent clusters to help avoid collapsed aggregated prototype representations.
Implication for Convergence
FedCC does not interfere with the DNN weight averaging step of FedAvg (used identically in FUSS), whose convergence under standard assumptions is supported by prior work mainly through empirical results [22]. Given that the centroid aggregation step in FedCC either:
  • 1.
    matches FedAvg when prototypes are aligned thus inherits convergence,
  • 2.
    or improves semantic consistency when FedAvg fails (without altering weight optimization),
we argue that FedCC introduces no instability. In fact, it enhances robustness against semantic collapse without modifying the underlying convergence properties of the global model.
The key technical role of FedCC can be summarized as follows: In cases where FedAvg performs well, FedCC matches its convergence behavior. However, in cases where FedAvg produces suboptimal class prototypes via naive centroid averaging, FedCC consistently improves performance by guiding local training toward more discriminative and stable semantic clusters.
@@ 5.1 Experimental Setup
We conduct experiments on the following benchmarks: Cityscapes [7], a 1/21 subsampled version of CocoStuff [1], and a private industrial dataset referred to as IPS (Industrial Pipeline Segmentation) [30].
5.1.1 Cityscapes
@@
We employ the Cityscapes dataset to simulate cross-silo federation: we reserve all samples from three cities for validation, and partition the remaining cities among clients. Depending on the federation size (3, 6, or 18 clients), each client is assigned images from 6, 3, or 1 city respectively, promoting spatial and semantic locality in client data.
Each federation client $k$ receives a disjoint subset $\mathcal{D}_{k}$ composed of images from geographically distinct cities, introducing domain distribution shifts. An example of the Cityscapes split for each federation simulation can be observed in Table 1. We also apply a five-crop on each image as a data augmentation technique, to better handle training times and small objects. It is important to note that while class distributions remain approximately i.i.d. across clients, the data still emulate a realistic federated learning setup in which each silo originates from a unique geographic location, thereby inducing domain and background variability among clients.
@@
Table 1: Cityscapes benchmark split for applied cross-silo federation scenarios
@@
5.1.2 CocoStuff
Although the CocoStuff dataset is not inherently designed for real-world federated learning scenarios, it is employed here to evaluate the robustness of our framework under extreme inter-client class heterogeneity. Each image is assigned a dominant semantic label based on the pixel-wise frequency of classes, after which a Dirichlet-based sampling strategy is applied to partition the dataset among clients. This procedure induces a controlled yet pronounced class imbalance and variability across clients.
We construct non-i.i.d. partitions of the CocoStuff dataset using a Dirichlet-based sampling strategy informed by semantic supercategories. The process involves assigning each image to a single coarse semantic class based on pixel frequency, followed by Dirichlet sampling to allocate images to clients in a skewed but controllable manner.